Q 3.27 Calculate the number of C, H, and O atoms in 1.50 g of glucose (C6H12O6), a sugar.


1
Expert's answer
2021-10-22T03:53:11-0400

v(c6h12o6)=1.5/180=0.008mol

m(C)=6*0.008*12=0.6g

m(H)=12*0.008*1=0.1g

m(O)=6*0.008*16=0.8g